OpenStudy (anonymous):

WILL GIVE MEDAL! Solve each equation for x. ax+7=5

OpenStudy (anonymous):

Theres also 1/ax-4=9 that i don't get

OpenStudy (anonymous):

ax+7=5 -7 -7 ax=-2 divide each side by a x=-2/a

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

do you understand how to do the second one?

OpenStudy (anonymous):

Oh no i don't

OpenStudy (anonymous):

ooops, x=1/13a

OpenStudy (anonymous):

It's not multiple choice, my teacher just wants us to try to solve it

OpenStudy (anonymous):

\[\frac{ 1 }{ ax }-4=9\]

OpenStudy (anonymous):

thats the equation right?

OpenStudy (anonymous):

It's with the x beside the 1/a

OpenStudy (anonymous):

]\[\frac{ 1 }{ a }x\]

OpenStudy (anonymous):

\[\frac{ 1 }{ a }x-4=9\]

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

alright, then its x=13a

OpenStudy (anonymous):

Okay thank you for taking your time to help me!

OpenStudy (anonymous):

no problem!